# Improved High-probability Convergence Guarantees of Decentralized SGD

### v2

@@ 5 Conclusion
In this paper, we revisit convergence in HP of a variant of DSGD under sub-Gaussian noise. We show that DSGD is guaranteed to converge in the HP sense under the same conditions on the cost as in the MSE sense, achieving order-optimal rates for both non-convex and strongly convex costs. Moreover, our results show DSGD achieves linear speed-up in the number of users in both cases. Compared to [dsgd-high-prob], wherein the authors study HP convergence of DSGD, we relax strong conditions like uniformly bounded gradients or asymptotically vanishing noise, while showing improved rates, with linear speed-up in the number of users and transient times matching, or strictly sharper than transient times stemming from corresponding MSE rates. Compared to works studying HP convergence of centralized SGD, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high, liu2024revisiting], we provide several novel results, establishing tight control on the MGF of the consensus gap for both non-convex and strongly convex costs in Lemmas 3 and 4, a more fine-grained bound on the MGF of an “almost decreasing” process in Lemma 5, as well as showing variance reduction benefits of decentralized algorithms (in the HP sense) in Lemma 2. Future work includes extending our results to costs satisfying the PŁ condition, incorporating the GT mechanism to remove bounded heterogeneity condition and considering noise with heavier tails, like sub-Weibull.
@@ Appendix B Useful Inequalities
Next, we state a useful result from [ran-improved].
Proposition 6 ([ran-improved], Lemma 21).
@@
$\prod_{k=a}^{b}\Big(1-\frac{c}{t+t_{0}}\Big)\leq\frac{(a+t_{0})^{c}}{(b+1+t_{0})^{c}}.$
The following result states some important consequences of conditions (A2) and (A3), see, e.g., [bertsekas-gradient, nesterov-lectures_on_cvxopt].
@@
Finally, we complete this section with the following important consequences of (A6), see, e.g., [nesterov-lectures_on_cvxopt].
@@ Appendix C Technical Results
    To prove the first property, we follow steps similar to those in [li2020high, Lemma 1]. Let $y_{i}=\frac{z^{t}_{i}}{\sigma_{i}}$ and note that $\mathbb{E}\left[\exp\left(\|y_{i}\|^{2}\right)\>|\>\mathcal{F}_{t}\right]\leq\exp(1)$, from (A3). Assume first that $v\in\mathbb{R}^{d}$ is such that $\|v\|\leq\frac{4}{3}$. Using the inequality $\exp(a)\leq a+\exp(\nicefrac{{9a^{2}}}{{16}})$, which holds for any $a\in\mathbb{R}$, we then have
@@
    Combining the previous result with Lemma 1 from [jin2019short], it can be readily seen that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{3d}}{\sqrt{n}}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$, we have
@@
    Using the equivalence between different conditions for scalar sub-Gaussian random variables, see, e.g., [vershynin_2018, Proposition 2.5.2], with $\|\overline{z}^{t}\|$ being the variable of interest, it can be readily verified that $\mathbb{E}\Big[\exp\Big(\frac{\|\overline{z}^{t}\|^{2}}{K^{2}}\Big)\Big]\leq\exp(1)$, where $K\leq\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$, implying that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian.
@@ Appendix D Analysis Setup
We now introduce some notation useful for the analysis of decentralized methods, see, e.g., [nedic-subgrad, jakovetic-unification, ran-improved]. Let ${\mathbf{x}}^{t}\coloneq col(x_{1}^{t},\ldots,x_{n}^{t})\in\mathbb{R}^{nd}$ denote the column vector stacking users’ local models. Using this notation, we can then represent the update rule (12) compactly as
@@
where $\mathbf{W}=W\otimes I_{d}\in\mathbb{R}^{nd\times nd}$, $\otimes$ denotes the Kronecker product and $I_{d}\in\mathbb{R}^{d\times d}$ denotes the $d$-dimensional identity matrix, while ${\mathbf{g}}^{t}\coloneqq col(g^{t}_{1},\ldots,g^{t}_{n})$. Next, define the matrix $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$, where $\mathbf{1}_{n}\in\mathbb{R}^{n}$ is the vector of all ones. The matrix $J$ represent the “ideal” consensus matrix, where all users can communicate with each other. [^13] This is equivalent to the client-server setup in terms of the update rule, as the server averages models from all clients in each iteration. The interaction between matrices $W$ and $J$ represents an important part of any decentralized algorithm and we now list some known properties, see, e.g., [jakovetic-unification, ran-improved] and references therein.
@@
    $\|W-J\|=\lambda$, where $\lambda\in[0,1)$ is the second largest singular value of $W$.
@@ Appendix E Proofs for Non-convex Costs
The proof follows similar steps as in [ran-improved, Lemma 3]. Starting from Proposition 7 and averaging across all costs $i\in[n]$, it readily follows that, for all $x,y\in\mathbb{R}^{d}$
@@
    as the last term, which does not attain a linear speed-up, is no longer of higher order. We note that linear speed-up in the MSE sense is also achieved under a known time horizon and fixed step-size, e.g., [pmlr-v119-koloskova20a, ran-improved].
@@ Appendix F Proofs for Strongly Convex Costs
In this section we prove Lemma 4 and Theorem 2. To do so, we follow a similar strategy to the one in, e.g., [conv-rates-dsgd-jakovetic, ran-improved], where it is shown that the sequence of iterates generated by DSGD is bounded in the MSE sense. However, to establish guarantees in the HP sense, we instead work with the MGF of the iterates and have the following important result.
@@
Let (A1)-(A4) and (A6) hold, let $a,t_{0},K>0$ and the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$, and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\Big\{6,\frac{288\overline{\sigma}^{2}K}{\mu^{2}},\frac{3456\overline{\sigma}^{2}\lambda^{2}K}{\mu^{2}(1-\lambda)},\frac{12\lambda L\sqrt{10}}{\mu(1-\lambda)}\Big\}$, then for $K_{t+1}=(t+t_{0}+2)K$ and any $\nu\leq\min\left\{1,\frac{\mu^{2}}{144\sigma^{2}K}\right\}$
@@
where $S_{k}=\alpha_{k}^{2}\lambda^{2}\Big(n\sigma^{2}+\frac{5\|\nabla\mathbf{f}^{\star}\|^{2}}{1-\lambda}\Big)$ and $D_{k}=\frac{5\alpha_{k}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{k}}{5}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(k+t_{0})^{6}}\Big)$.
@@
We remark that, similarly to the proof of Theorem 1, one can analyze the strongly convex case with a fixed step-size, resulting in the dependence on some terms, e.g., optimality and iterate gaps $\Delta_{f}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}}^{\star}\|$, decaying exponentially fast, i.e., $\mathcal{O}\big((\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}+\Delta_{f})e^{-CT}\big)$, for some $C>0$, as shown in, e.g., [pmlr-v119-koloskova20a] for MSE guarantees in decentralized, or [liu2024revisiting] for HP guarantees in centralized settings. For simplicity, we omit this analysis.
@@ Appendix G Numerical Experiments
Methodology. We consider an instance of (2), with $f_{i}(x)=\frac{1}{2}x^{\top}A_{i}x+b_{i}^{\top}x$, where $A_{i}\in\mathbb{R}^{d\times d}$ is positive semi-definite, making each $f_{i}$ strongly convex, with unbounded gradients. We consider an undirected communication network $G=(V,E)$, corresponding to a randomly generated Erdős–Rényi graph with connectivity parameter $p=0.8$, while the weight matrix $W\in\mathbb{R}^{n\times n}$ is computed using the Metropolis-Hastings weight scheme, e.g., [XIAO200465]. When queried by user $i\in[n]$ in iteration $t\geq 1$, the $\mathcal{SFO}$ returns a noisy gradient of $f_{i}$ evaluated at $x^{t}_{i}$, i.e., $g^{t}_{i}=A_{i}x^{t}_{i}+b_{i}+z^{t}_{i}$, where $z^{t}_{i}\in\mathbb{R}^{d}$ is a zero-mean Gaussian random vector, making the noise consistent with assumption (A4). We use the DSGD method outlined in Algorithm 1, with the time-varying step-size schedule $\alpha_{t}=\frac{1}{t+1}$ and shared initializatio $x^{1}_{i}=0$, for all $i\in[n]$. We are interested in testing the following two facets of our theory.
@@
Exponentially decaying tails. To verify that the (empirical) tail probability decays at an exponential rate, we consider a fixed network of $n=30$ users. For local costs, each matrix $A_{i}$ is generated using python’s sklearn library function make_sparse_spd_matrix, with dimension $d=50$ and value $\texttt{alpha}=0.9$, while vectors $b_{i}$ are drawn from a multivariate normal distribution $\mathcal{N}(\mathbf{0}_{d},\sigma_{i}^{2}I_{d})$, where $\sigma_{i}^{2}=1$ for $i\in\{1,\ldots,10\}$, $\sigma_{i}^{2}=2$ for $i\in\{11,\ldots,20\}$, and $\sigma^{2}_{i}=4$ for $i\in\{21,\ldots,30\}$, ensuring that the data is heterogeneous across users. We run DSGD for $T=10000$ iterations and repeat across $R=5000$ runs. To compute the empirical probability, in each iteration we draw $S=3000$ Monte-Carlo samples and consider threshold values $\varepsilon=\big\{10^{-2},10^{-3},10^{-4}\big\}$. The results are presented in Figure 1, where the left plot shows the MSE behaviour, while the right plot shows the tail probability behaviour. We can see that the empirical tail probability decays exponentially fast for all values of $\varepsilon$, as predicted by our theory. Note that for the threshold value $\varepsilon=10^{-4}$, the tail probability starts decaying exponentially after approximately $t=6000$ iterations, which is consistent with the MSE behaviour on the left plot, where we can see that it takes DSGD around the same number of iterations to reach the average accuracy $\mathbb{E}_{n}^{t}=10^{-4}$.
@@
Linear speed-up. To verify that linear speed-up in the number of users is achieved, we consider three networks, with $n=\{25,50,100\}$ users. In all three cases we use Erdős–Rényi communications graphs which satisfy $\lambda=\|W-J\|\approx 0.6$. To ensure that the effect of heterogeneity is consistent across different networks, we fix the matrices $A_{i}$, i.e., $A_{i}\equiv A=S+I_{d}$, where $S\in\mathbb{R}^{d\times d}$ is generated using sklearn’s make_sparse_spd_matrix, with $d=50$ and $\texttt{alpha}=0.9$, while vectors $b_{i}$ are given by $b_{i}=\beta_{i}\mathbf{1}_{d}$, where $\beta_{i}\in\{-2,-1,0,1,3\}$ are selected uniformly at random and in equal proportion across users (i.e., one fifth of users has $\beta_{i}=-2$, one fifth $\beta_{i}=-1$, etc). Generating the networks and costs in this manner ensures that both network connectivity and user heterogeneity are constant across all three network settings, allowing us to properly capture the effect of linear speed-up. We run DSGD for $T=3000$ iterations and repeat across $R=1000$ runs. To compute the empirical probability, in each iteration we draw $S=600$ Monte-Carlo samples, with thresholds $\varepsilon=\big\{10^{-2},10^{-3},10^{-4}\big\}$. The results are presented in Figure 2, where the plots left to right and top to bottom show the MSE and tail probability behaviours for different values of $\varepsilon$, respectively. We can again see that the empirical tail probability decays exponentially fast for all values of $\varepsilon$, with the decay being consistently faster for larger number of users, demonstrating linear speed-up in the HP sense. Finally, note that in the lower right plot (i.e., threshold $\varepsilon=10^{-4}$) the network with $n=25$ users has a constant tail probability $\mathbb{P}^{t}_{n,\varepsilon}=1$, while for the network with $n=50$ the tail probability is slowly starting to decrease, which is again consistent with the MSE behaviour on the upper left plot, as the accuracy $\mathbb{E}^{t}_{n}=10^{-4}$ for these two networks is not reached (on average) in the allocated number of iterations.
@@ Appendix H Deriving the Transient Time
In this section we provide details on the transient time resulting from our HP bounds. Similarly to [unified-refined], we focus on the dependence on network connectivity and the number of users, ignoring other problem related constants. First, consider non-convex costs. Recall that the bound in Theorem 1 is of the form
@@
It is now evident that, for any $T\geq\max\Big\{\frac{n}{(1-\lambda)^{2}},\frac{n^{3}}{(1-\lambda)^{4}}\Big\}=\frac{n^{3}}{(1-\lambda)^{4}}$, the bound in (46) becomes $\mathcal{O}\Big(\frac{1}{\sqrt{nT}}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$. This matches the transient time of DSGD obtained from the MSE rate in [pmlr-v119-koloskova20a], see, e.g., [unified-refined, Table I].
@@
where $t_{0}$ depends on the network connectivity via the expression $t_{0}=\Omega\Big(\max\Big\{\frac{\lambda}{1-\lambda},\frac{\lambda^{2}}{1-\lambda},\frac{1}{1-\lambda}\Big\}\Big)$. Since we are again interested in the worst-case transient time, it follows that $t_{0}=\Omega\Big(\frac{\lambda}{1-\lambda}\Big)$, so that the bound in (H) becomes
@@
It can be seen that, for any $T\geq\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}},\frac{\sqrt{n}}{1-\lambda},\frac{1}{(1-\lambda)^{\frac{3}{2}}},\frac{1}{(1-\lambda)^{\frac{8}{7}}}\Big\}=\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}$, the bound in (H) becomes $\mathcal{O}\Big(\frac{\log(T+t_{0})}{n(T+t_{0})}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$. Interestingly, this is strictly sharper than the transient time of DSGD obtained from the MSE rate in [pmlr-v119-koloskova20a], which is of the order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see, e.g., [unified-refined, Table II]. As such, our HP analysis provides either matching or strictly sharper transient times for DSGD compared to the transient times stemming from the corresponding MSE results, further highlighting the tightness of our analysis.

### v3

@@ 5 Conclusion
We studied convergence in HP of a variant of DSGD under light-tailed noise, showing that it is guaranteed to converge in the HP sense under the same conditions on the cost as in the MSE sense, achieving order-optimal rates and linear speed-up for both non-convex and strongly convex costs. Compared to existing works, we relax strong assumptions like uniformly bounded gradients and asymptotically vanishing noise, while simultaneously showing improved rates and sharper transient times than the ones obtained from MSE rates. Future work includes extending our results to costs satisfying the PŁ condition, incorporating bias-correction mechanisms like GT and exact diffusion to remove bounded heterogeneity, and considering heavy-tailed noise.
@@ Appendix B Useful Inequalities
Next, we state a useful result from [75].
Proposition 6 ([75], Lemma 21).
@@
$\prod_{k=a}^{b}\Big(1-\frac{c}{k+t_{0}}\Big)\leq\frac{(a+t_{0})^{c}}{(b+1+t_{0})^{c}}.$
The following result states some important consequences of conditions (A2) and (A3), see, e.g., [9, 55].
@@
Finally, we complete this section with the following important consequences of (A6), see, e.g., [55].
@@ Appendix C Technical Results
    To prove the first property, we follow steps similar to those from Lemma 1 in [37]. Let $y_{i}=\frac{z^{t}_{i}}{\sigma_{i}}$ and note that $\mathbb{E}\left[\exp\left(\|y_{i}\|^{2}\right)\>|\>\mathcal{F}_{t}\right]\leq\exp(1)$, from (A3). Assume first that $v\in\mathbb{R}^{d}$ is such that $\|v\|\leq\frac{4}{3}$. Using the inequality $\exp(a)\leq a+\exp(\nicefrac{{9a^{2}}}{{16}})$, which holds for any $a\in\mathbb{R}$, we then have
@@
    Combining the previous result with Lemma 1 from [27], it can be readily seen that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{3d}}{\sqrt{n}}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$, we have
@@
    Using the equivalence between different conditions for scalar sub-Gaussian random variables, see, e.g., Proposition 2.5.2 in [67], with $\|\overline{z}^{t}\|$ being the variable of interest, it can be readily verified that $\mathbb{E}\Big[\exp\Big(\frac{\|\overline{z}^{t}\|^{2}}{K^{2}}\Big)\Big]\leq\exp(1)$, where $K\leq\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$, implying that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian.
@@ Appendix D Analysis Setup
We now introduce some notation useful for the analysis of decentralized methods, see, e.g., [53, 26, 75]. Let ${\mathbf{x}}^{t}\coloneq col(x_{1}^{t},\ldots,x_{n}^{t})\in\mathbb{R}^{nd}$ denote the column vector stacking users’ local models. Using this notation, we can then represent the update rule (14) compactly as
@@
where $\mathbf{W}=W\otimes I_{d}\in\mathbb{R}^{nd\times nd}$, $\otimes$ denotes the Kronecker product and $I_{d}\in\mathbb{R}^{d\times d}$ denotes the $d$-dimensional identity matrix, while ${\mathbf{g}}^{t}\coloneqq col(g^{t}_{1},\ldots,g^{t}_{n})$. Next, define the matrix $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$, where $\mathbf{1}_{n}\in\mathbb{R}^{n}$ is the vector of all ones. The matrix $J$ represent the “ideal” consensus matrix, where all users can communicate with each other. [^10] This is equivalent to the client-server setup in terms of the update rule, as the server averages models from all clients in each iteration. The interaction between matrices $W$ and $J$ represents an important part of any decentralized algorithm and we now list some known properties, see, e.g., [26, 75] and references therein.
@@
    $\lambda\coloneqq\|W-J\|\in[0,1)$.
@@ Appendix E Proofs for Non-convex Costs
The proof follows similar steps as in Lemma 3 from [75]. Starting from Proposition 7 and averaging across all costs $i\in[n]$, it readily follows that, for all $x,y\in\mathbb{R}^{d}$
@@
    as the last term, which does not attain a linear speed-up, is no longer of higher order. We note that linear speed-up in the MSE sense is also achieved under a known time horizon and fixed step-size, e.g., [33, 75].
@@ Appendix F Proofs for Strongly Convex Costs
In this section we prove Lemma 4 and Theorem 2. To do so, we follow a similar strategy to the one in, e.g., [24, 75], where it is shown that the sequence of iterates generated by DSGD is bounded in the MSE sense. However, to establish guarantees in the HP sense, we instead work with the MGF of the iterates and have the following important result.
@@
Let (A1)-(A4) and (A6) hold, let $a,t_{0},K>0$ and the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$, and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\Big\{6,\frac{288\overline{\sigma}^{2}K}{\mu^{2}},\frac{3456\overline{\sigma}^{2}\lambda^{2}K}{\mu^{2}(1-\lambda)},\frac{12\lambda L\sqrt{10}}{\mu(1-\lambda)}\Big\}$, then for $K_{t+1}=(t+t_{0}+2)K$ and any $\nu\leq\min\left\{1,\frac{\mu^{2}}{144\sigma^{2}K}\right\}$, we have
@@
We remark that, similarly to the proof of Theorem 1, one can analyze the strongly convex case with a fixed step-size, resulting in the dependence on some terms, e.g., optimality and iterate gaps $\Delta_{f}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}}^{\star}\|$, decaying exponentially fast, i.e., $\mathcal{O}\big((\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}+\Delta_{f})e^{-CT}\big)$, for some $C>0$, as shown in, e.g., [33] for MSE guarantees in decentralized, or [41] for HP guarantees in centralized settings. For simplicity, we omit this analysis.
@@ Appendix G Numerical Experiments
Methodology.
We consider an instance of (2), with $f_{i}(x)=\frac{1}{2}x^{\top}A_{i}x+b_{i}^{\top}x$, where $A_{i}\in\mathbb{R}^{d\times d}$ is positive semi-definite, making each $f_{i}$ strongly convex, with unbounded gradients. We consider an undirected communication network $G=(V,E)$, corresponding to a randomly generated Erdős–Rényi graph with connectivity parameter $p=0.8$, while the weight matrix $W\in\mathbb{R}^{n\times n}$ is computed using the Metropolis-Hastings weight scheme, e.g., [73]. When queried by user $i\in[n]$ in iteration $t\geq 1$, the $\mathcal{SFO}$ returns a noisy gradient of $f_{i}$ evaluated at $x^{t}_{i}$, i.e., $g^{t}_{i}=A_{i}x^{t}_{i}+b_{i}+z^{t}_{i}$, where $z^{t}_{i}\in\mathbb{R}^{d}$ is a zero-mean Gaussian random vector, making the noise consistent with assumption (A4). We use the DSGD method outlined in Algorithm 1, with the time-varying step-size schedule $\alpha_{t}=\frac{1}{t+1}$ and shared initializatio $x^{1}_{i}=0$, for all $i\in[n]$. We are interested in testing the following two facets of our theory.
@@
Exponentially decaying tails.
To verify that the (empirical) tail probability decays at an exponential rate, we consider a fixed network of $n=30$ users. For local costs, each matrix $A_{i}$ is generated using python’s sklearn library function make_sparse_spd_matrix, with dimension $d=50$ and value $\texttt{alpha}=0.9$, while vectors $b_{i}$ are drawn from a multivariate normal distribution $\mathcal{N}(\mathbf{0}_{d},\sigma_{i}^{2}I_{d})$, where $\sigma_{i}^{2}=1$ for $i\in\{1,\ldots,10\}$, $\sigma_{i}^{2}=2$ for $i\in\{11,\ldots,20\}$, and $\sigma^{2}_{i}=4$ for $i\in\{21,\ldots,30\}$, ensuring that the data is heterogeneous across users. We run DSGD for $T=10000$ iterations and repeat across $R=5000$ runs. To compute the empirical probability, in each iteration we draw $S=3000$ Monte-Carlo samples and consider threshold values $\varepsilon=\big\{10^{-2},10^{-3},10^{-4}\big\}$. The results are presented in Figure 1, where the left plot shows the MSE behaviour, while the right plot shows the tail probability behaviour. We can see that the empirical tail probability decays exponentially fast for all values of $\varepsilon$, as predicted by our theory. Note that for the threshold value $\varepsilon=10^{-4}$, the tail probability starts decaying exponentially after approximately $t=6000$ iterations, which is consistent with the MSE behaviour on the left plot, where we can see that it takes DSGD around the same number of iterations to reach the average accuracy $\mathbb{E}_{n}^{t}=10^{-4}$.
@@
Linear speed-up.
To verify that linear speed-up in the number of users is achieved, we consider three networks, with $n=\{25,50,100\}$ users. In all three cases, we use Erdős–Rényi communications graphs with Metropolis-Hastings weights that satisfy $\lambda=\|W-J\|\approx 0.6$. To ensure that the effect of heterogeneity is consistent across different networks, we fix the matrices $A_{i}$, i.e., $A_{i}\equiv A=S+I_{d}$, where $S\in\mathbb{R}^{d\times d}$ is generated using sklearn’s make_sparse_spd_matrix, with $d=50$ and $\texttt{alpha}=0.9$, while vectors $b_{i}$ are given by $b_{i}=\beta_{i}\mathbf{1}_{d}$, where $\beta_{i}\in\{-2,-1,0,1,3\}$ are selected uniformly at random and in equal proportion across users (i.e., one fifth of users has $\beta_{i}=-2$, one fifth has $\beta_{i}=-1$, etc). Generating the networks and costs in this manner ensures that both network connectivity and user heterogeneity are constant across all three network settings, allowing us to properly capture the effect of linear speed-up. We run DSGD for $T=3000$ iterations and repeat across $R=1000$ runs. To compute the empirical probability, in each iteration we draw $S=600$ Monte-Carlo samples, with thresholds $\varepsilon=\big\{10^{-2},10^{-3},10^{-4}\big\}$. The results are presented in Figure 2, where the plots left to right and top to bottom respectively show the MSE and tail probability behaviours for different values of $\varepsilon$. We can again see that the empirical tail probability decays exponentially fast for all values of $\varepsilon$, with the decay being consistently faster for larger number of users, demonstrating linear speed-up in the HP sense. Finally, note that in the lower right plot (i.e., threshold $\varepsilon=10^{-4}$) the network with $n=25$ users has a constant tail probability $\mathbb{P}^{t}_{n,\varepsilon}=1$, while for the network with $n=50$ the tail probability is slowly starting to decrease, which is again consistent with the MSE behaviour on the upper left plot, as the accuracy $\mathbb{E}^{t}_{n}=10^{-4}$ for these two networks is not reached (on average) in the allocated number of iterations.
@@ Appendix H Deriving the Transient Time
In this section we provide details on the transient time resulting from our HP bounds. Similarly to [1], we focus on the dependence on network connectivity and the number of users, ignoring other problem related constants. First, consider non-convex costs. Recall that the bound in Theorem 1 is of the form
@@
It is now evident that, for any $T\geq\max\Big\{\frac{n}{(1-\lambda)^{2}},\frac{n^{3}}{(1-\lambda)^{4}}\Big\}=\frac{n^{3}}{(1-\lambda)^{4}}$, the bound in (48) becomes $\mathcal{O}\Big(\frac{1}{\sqrt{nT}}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$. This matches the transient time of DSGD obtained from the MSE rate in [33], see, e.g., Table I in [1].
@@
where $t_{0}$ depends on the network connectivity via the expression $t_{0}=\Omega\Big(\max\Big\{\frac{\lambda}{1-\lambda},\frac{\lambda^{2}}{1-\lambda},\frac{1}{1-\lambda}\Big\}\Big)$. Since we are again interested in the worst-case transient time, it follows that $t_{0}=\Omega\Big(\frac{1}{1-\lambda}\Big)$, so that the bound in (H) becomes
@@
It can be seen that, for any $T\geq\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}},\frac{\sqrt{n}}{1-\lambda},\frac{1}{(1-\lambda)^{\frac{3}{2}}},\frac{1}{(1-\lambda)^{\frac{8}{7}}}\Big\}=\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}$, the bound in (H) becomes $\mathcal{O}\Big(\frac{\log(T+t_{0})}{n(T+t_{0})}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$. Interestingly, this is strictly sharper than the transient time of DSGD obtained from the MSE rate in [33], which is of the order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see, e.g., Table II in [1]. As such, our HP analysis provides either matching or strictly sharper transient times for DSGD compared to the transient times stemming from the corresponding MSE results, further highlighting the tightness of our analysis.